A Company reports revenues, for financial reporting purpose, equal to $100,000, which includes $10,000 of revenue uncollected. Therefore, for tax purpose, the company reports revenue equal to $90,000 ($100,000-$10,000). The company also reports expenses for both financial reporting purpose and tax purpose equal to $60,000.

Assuming statutory tax rate is 35%.

a. How much is "income taxes paid"? 10,500

c. The amount of the deferred tax asset/liability created is ____________

d. Income tax expense is calculated as _____________________

e. Effective tax rate is equal to (use percentage) __________________

Please explain how to get the answers (or post work) for c,d,e so I can understand. Thank you.
Answer #1

a

Income tax paid = taxable income X tax rate = 30,000 X 35% = 10,500

c

Deferred tax liability = revenue not taxed now X tax rate = 10,000 X 35% = 3,500

d

Income tax expense = (Book income - permanent differences) X tax rate = (40,000 - 0) X 35% = 14,000

e

Effective tax rate = Income tax expense/ book income = 14,000/40,000 = 35%